OpenStudy (anonymous):

i dont seem to be able to use the divide sign in LaTeX... what could be the problem??

OpenStudy (anonymous):

sometimes when i use the 'over' command, i dont see the horizontal bar that separates the numerator from denominator.

OpenStudy (mattfeury):

I'm not too familiar with latex, but in the equation dialog I usually use: frac{num}{denom} rendered like so: \[\frac{\pi}{2}\]

OpenStudy (shadowfiend):

Yes, I've noticed this bug. It has to do with our math typesetting library; we'll be updating to its latest version very soon and that should hopefully help.

OpenStudy (shadowfiend):

And to be clear -- both \frac and \over are susceptible to it, unfortunately. Also, I've been using LaTeX for a while and this is the first time I've heard of \over. < shakes fist > :p